# Sobolev inequality and its applications to nonlinear PDE on noncommutative Euclidean spaces

### v1

@@ 1. Introduction
There are several proofs of the Sobolev inequality in One of them is by using the heat kernel estimate [39]. Indeed, N. Varopoulos obtained an equivalence between a heat kernel estimate and the Sobolev inequality in an abstract settings in 1985 [38]. He proved that the Sobolev inequality is equivalent to the heat kernel estimate as well as the Nash, and Moser inequalities, for possibly different constants. On the other hand, the Nash inequality is also equivalent to the logarithmic Sobolev inequality. The Sobolev inequality was proved in [37, Theorem 3.11 and Corollary 3.19] by using the heat kernel estimate. Moreover, the author obtained the noncommutative analogue of the Varopoulos’s theorem even a more general case which includes noncommutative Euclidean spaces. On the other hand, there is another way to prove the Sobolev inequality which hinges upon the Young inequality for weak type spaces and the Hardy-Littlwood-Sobolev inequality. Recently the Sobolev inequality was also studied in the quantum phase space in [18] by this way. We think that the quantum phase space is equivalent to the particular case of our noncommutative Euclidean space when In this work, we study general case which covers some results in [18]. Therefore, we first focused to obtain the Hardy-Littlwood-Sobolev and Young inequality for weak type spaces in the noncommutative Euclidean space. Our proof is based on the Young inequality in [21] and the interpolation in the classical Lorentz spaces as well as a proper extension of the definition of convolutions in the class of distributions. Moreover, we prove that the equivalence between the Nash inequality the logarithmic Sobolev inequality remains true even in the noncommutative Euclidean space by completing the list of the noncommutative Varopoulos’s theorem in [37, Theorem 4.30]. As a consequence of the noncommutative Sobolev and Hölder inequalities we obtain a version of the Gagliardo–Nirenberg inequality which allows us to show its global well-posedness of nonlinear damped wave equations for the sub-Laplacian in the noncommutative Euclidean space. Nonlinear PDEs in the noncommutative Euclidean space were studied very recently in [21], where the author obtained local (some global) well-posedness in time of Allen-Cahn, Schrödinger and incompressible Navier-Stokes equations. The main difference in this work is that we study damped wave equations for the sub-Laplacian and show its global in time well posedness. Moreover, our approach is completely different from that of [21]. However, our approach is also applicable at least to obtain a global in time well posedness of the nonlinear Schrödinger equation which was studied in [21]. In general, there are some technical difficulties in noncommutative setting including defining even nonlinear operator functions. The other difficulties in the noncommutative Euclidean spaces were explained in [21]. The main tool to show that is the Gagliardo–Nirenberg type inequality and proper Banach fixed point theorem. The idea comes from the paper [29], where the authors studied similar problems in the context of Heisenberg and graded Lie groups. At the end, as in the classical case, we show a direct application of the Nash inequality to compute the time decay for solutions of the heat equation in the noncommutative setting.
@@ 2.6. Convolution on
[37, Theorem 3.11 and Corollary 3.19] Let Then the following inequalities are equivalent,
@@ 4. Equivalence of various inequalities
4.1. The equivalence of Nash inequality and logarithmic Sobolev inequality
@@
(Logarithmic Sobolev inequality). Let be such that Then for any such that we have the fractional logarithmic Sobolev inequality
@@
Let Then for nay such that the following inequalities are equivalent:
@@
differentiating the function with respect to and we have
@@
Let Then for any such that the following inequalities are equivalent:
@@ 5.2. Linear damped wave equation on noncommutative Euclidean spaces
The Sobolev space with associated to operator is defined as
@@ 5.3. Semilinear damped wave equations on noncommutative Euclidean spaces
The absolute value function is Lipschitz continuous on all -spaces for as demonstrated in the self-adjoint case in [5]. For the case of normal operators, refer to the paper by [17]. This result enables the application of the absolute value function to your operators and allowing you to focus exclusively on the case of positive operators.

### v5

@@ 1. Introduction
There are several proofs of the Sobolev inequality in One of them is by using the heat kernel estimate [40]. Indeed, N. Varopoulos obtained an equivalence between a heat kernel estimate and the Sobolev inequality in an abstract settings in 1985 [39]. He proved that the Sobolev inequality is equivalent to the heat kernel estimate as well as the Nash, and Moser inequalities, for possibly different constants. On the other hand, the Nash inequality is also equivalent to the logarithmic Sobolev inequality. The Sobolev inequality was proved in [38, Theorem 3.11 and Corollary 3.19] by using the heat kernel estimate. Moreover, the author obtained the noncommutative analogue of the Varopoulos’s theorem even a more general case which includes noncommutative Euclidean spaces. On the other hand, there is another way to prove the Sobolev inequality which hinges upon the Young inequality for weak type spaces and the Hardy-Littlwood-Sobolev inequality. Recently the Sobolev inequality was also studied in the quantum phase space in [18] by this way. We think that the quantum phase space is equivalent to the particular case of our noncommutative Euclidean space when In this work, we study general case which covers some results in [18]. Therefore, we first focused to obtain the Hardy-Littlwood-Sobolev and Young inequality for weak type spaces in the noncommutative Euclidean space. Our proof is based on the Young inequality in [21] and the interpolation in the classical Lorentz spaces as well as a proper extension of the definition of convolutions in the class of distributions. Moreover, we prove that the equivalence between the Nash inequality the logarithmic Sobolev inequality remains true even in the noncommutative Euclidean space by completing the list of the noncommutative Varopoulos’s theorem in [38, Theorem 4.30] (see also [34] for the equivalence between -ultracontractivity and logarithmic Sobolev inequality for some in noncommutative setting). As a consequence of the noncommutative Sobolev and Hölder inequalities we obtain a version of the Gagliardo–Nirenberg inequality which allows us to show its global well-posedness of nonlinear damped wave equations for the sub-Laplacian in the noncommutative Euclidean space. Nonlinear PDEs in the noncommutative Euclidean space were studied very recently in [21], where the author obtained local (some global) well-posedness in time of Allen-Cahn, Schrödinger and incompressible Navier-Stokes equations. The main difference in this work is that we study damped wave equations for the sub-Laplacian and show its global in time well posedness. Moreover, our approach is completely different from that of [21]. However, our approach is also applicable at least to obtain a global in time well posedness of the nonlinear Schrödinger equation which was studied in [21]. In general, there are some technical difficulties in noncommutative setting including defining even nonlinear operator functions. The other difficulties in the noncommutative Euclidean spaces were explained in [21]. The main tool to show that is the Gagliardo–Nirenberg type inequality and proper Banach fixed point theorem. The idea comes from the paper [29], where the authors studied similar problems in the context of Heisenberg and graded Lie groups. At the end, as in the classical case, we show a direct application of the Nash inequality to compute the time decay for solutions of the heat equation in the noncommutative setting.
@@ 2.6. Convolution on
[38, Theorem 3.11 and Corollary 3.19] Let Then the following inequalities are equivalent,
@@ 4. Equivalence of various inequalities
4.1. Equivalence of the Nash and the logarithmic Sobolev inequalities
@@
(Logarithmic Sobolev inequality). Let be such that Then for any we have the fractional logarithmic Sobolev inequality
@@
Let Then for any the following inequalities are equivalent:
@@
differentiating the function with respect to (see [34, Lemma 4.2]), and we have
@@
Let Then for any the following inequalities are equivalent:
@@ 5.2. Linear damped wave equation on noncommutative Euclidean spaces
The Sobolev space with associated to the operator is defined as
@@ 5.3. Semilinear damped wave equations on noncommutative Euclidean spaces
The absolute value function is Lipschitz continuous on all -spaces for as demonstrated in the self-adjoint case in [5]. For the case of normal operators, we refer the reader to the paper [17]. These results enable the application of the absolute value function to operators and allowing us to focus exclusively on the case of positive operators.
@@ 5.4. Application of the Nash inequality to the heat equation
5.5. Conflict of interest
We can conceive of no conflict of interest in the publication of this paper. The work has not been published previously and it has not been submitted for publication elsewhere.
5.6. Data availability
No new data were created or analysed during this study. Data sharing is not applicable to this article